Suppose that AD, BC, AC, and BD are line segments with line AD parallel to line BC. If AD=3, BC=1, and the distance from AD to BC is 5, then what is the sum of the areas of the two triangles formed?
Answer by greenestamps(13198)



The two triangles are similar, because BC and AD are parallel.
The ratio of similarity is BC:AD = 1:3.
So the ratio of the altitudes of the two triangles is 1:3.
The (perpendicular) distance between BC and AD is 5.
Therefore the altitude of the small triangle is 1/4 of 5, or 5/4.
Then the area of the small triangle is one-half base times height = (1/2)(1)(5/4) = 5/8.
Since the ratio of similarity between the two triangles is 1:3, the ratio of their areas is 1^2:3^2, or 1:9.
So the area of the large triangle is 9*(5/8) = 45/8.
And so the sum of the areas of the two triangles is 5/8+45/8 = 50/8 = 25/4.